OpenStudy (anonymous):

In a certain county, the number of Charter Schools is eight less than twice the number of Alternative Schools. We know that there are 36 Charter Schools in the county. How many Alternative Schools are in the county?

OpenStudy (anonymous):

i got it. do u know how to do a algebraic expression????

OpenStudy (anonymous):

well its 8-2x then you divide by 2 on both sides so.... 8/2-2x/2= answer i got is 4 Hope this helps? (:

OpenStudy (texaschic101):

c = 2a - 8 c = 36 36 = 2a - 8 36 + 8 = 2a 44 = 2a 22 = a check.. 36 = 2(22) - 8 36 = 44 - 8 36 = 36 (correct) there are 22 alternative schools.